OpenStudy (anonymous):

What is the exact area of a circle with a diameter of m? A. m² B. m² C. m² D. m²

OpenStudy (anonymous):

area of a circle = (pi)(r^2) A=(pi)(m^2) i don't know why all your multiple choice answers are m^2, but my answer is correct.

OpenStudy (anonymous):

oops sorry i thought you meant radius was m. in this case, radius=diameter/2. so r=m/2. so A=(pi)(radius^2) A=(pi)(m/2)^2 A=(pi)(m^2/4) A=[(pi)(m^2)]/4